# Diffraction by a Right-Angle Penetrable Wedge: Closed-Form Solution for $\nu=\sqrt{2}$

### v2

@@ Abstract
MSC 2020: 35J05, 78A45, 33E05, 45E10.
Keywords: penetrable wedge diffraction; impedance-matched penetrable wedge; Sommerfeld integrals; Wiener–Hopf; Riemann–Hilbert; elliptic functions; lemniscatic curve; Weierstrass zeta-function.
@@
The author used OpenAI’s ChatGPT (GPT-5.2 Pro) as a writing and editing aid for exposition, LaTeX typesetting, and consistency checks. All mathematical derivations, claims, and proofs remain the responsibility of the author.
@@
    §11 proves that $Q_{\mathrm{scat}}$ is analytic at the incident spectral point $\zeta=\zeta_{i}$ (limiting absorption), by the explicit exclusion of the incident label from the scattered pole set.
@@
  • The complex spectral variable is $\zeta\in\mathbb{C}$ (Sommerfeld strip). The incident spectral pole is $\zeta_{i}=\theta_{i}+i\varepsilon$ with $\varepsilon>0$ (limiting absorption), and the physical limit is $\varepsilon\to 0^{+}$.
  • We use the Sommerfeld integration parameter $z$ and the associated variable $t=e^{iz}$.
  • The lemniscatic curve is $\Sigma_{\mathrm{lem}}:=\{(t,Y):\;Y^{2}=2(t^{4}+1)\}$. The physical sheet $\Omega_{+}^{\mathrm{phys}}\subset\Sigma_{\mathrm{lem}}$ is characterized by $|t|<1$ and $Y\to-\sqrt{2}$ as $t\to 0$.
  • $\zeta_{W}(u)$, $\wp(u)$, and $\wp^{\prime}(u)$ denote the Weierstrass zeta and elliptic functions with invariants $(g_{2},g_{3})=(4,0)$ (square lattice $\tau=i$); the subscript distinguishes the Weierstrass zeta function from the spectral variable $\zeta$.
@@ 1 Introduction and setup
Scope. The analysis and the resulting closed form apply only to the special configuration $(\theta_{w},\nu,\rho)=(\pi/4,\sqrt{2},1)$. We do not claim an explicit closed-form solution for general penetrable wedges (arbitrary contrast and wedge angle), for which the standard spectral reductions lead to matrix/generalized Wiener–Hopf or multi-variable boundary-value problems; see [4, 7, 9, 10, 12, 11, 13].
@@
$(\Delta+k_{0}^{2})u_{0}=0\quad\text{in }\{|\theta|>\theta_{w}\},\qquad(\Delta+k_{1}^{2})u_{1}=0\quad\text{in }\{|\theta|<\theta_{w}\},$
@@
$u_{\mathrm{inc}}(r,\theta)=\exp\!\bigl(ik_{0}r\cos(\theta-\theta_{i})\bigr),$
@@
$u_{0}(r,\pm\theta_{w})=u_{1}(r,\pm\theta_{w}),\qquad\partial_{\theta}u_{0}(r,\pm\theta_{w})=\partial_{\theta}u_{1}(r,\pm\theta_{w}),\qquad r>0.$
@@
$\zeta_{i}:=\theta_{i}+i\varepsilon,\qquad\varepsilon>0.$ (1)
@@ 1.1 Sommerfeld representation and spectral split
$u^{(0)}(r,\theta)=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{0}r\cos z}\Bigl(Q(\theta+z)-Q(\theta-z)\Bigr)\,dz,$ (2)
@@
$u^{(1)}(r,\theta)=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{1}r\cos z}\Bigl(S(\theta+z)-S(\theta-z)\Bigr)\,dz,\qquad|\theta|<\theta_{w},$ (3)
@@ 1.3 Transmission conditions in spectral form
with the branch determined by $w_{b}(z)\sim z+i\log\nu$ as $\Im z\to+\infty$. Write $w_{b}^{\prime}(z)=dw_{b}/dz$.
@@
$U(r,\theta):=\frac{1}{2\pi i}\int_{\gamma}e^{ikr\cos\zeta}\,H(\zeta)\,d\zeta.$
@@
Assume $Q$ and $S$ are analytic in a common Sommerfeld strip and have sufficient decay so that integration by parts in $z$ produces no boundary terms. Fix a face $\theta=\theta_{b}$ and let $w=w_{b}(z)$ be as in (7). Then the impedance-matched transmission conditions
$u_{0}(r,\theta_{b})=u_{1}(r,\theta_{b}),\qquad\partial_{\theta}u_{0}(r,\theta_{b})=\partial_{\theta}u_{1}(r,\theta_{b}),\qquad r>0,$
are equivalent to the pointwise spectral relation
$\begin{pmatrix}S(\theta_{b}+z)\\[2.0pt] S(\theta_{b}-z)\end{pmatrix}=\frac{1}{2}\begin{pmatrix}1+w_{b}^{\prime}(z)&1-w_{b}^{\prime}(z)\\[2.0pt] 1-w_{b}^{\prime}(z)&1+w_{b}^{\prime}(z)\end{pmatrix}\begin{pmatrix}Q(\theta_{b}+w_{b}(z))\\[2.0pt] Q(\theta_{b}-w_{b}(z))\end{pmatrix}$. (8)
@@
We record the standard argument, included here to make the reduction from the transmission problem explicit. Evaluate (2) and its $\theta$-derivative at $\theta=\theta_{b}$. Differentiating under the integral sign gives
$\partial_{\theta}u_{0}(r,\theta_{b})=\frac{1}{2\pi i}\int_{\gamma}e^{ik_{0}r\cos z}\bigl[Q^{\prime}(\theta_{b}+z)-Q^{\prime}(\theta_{b}-z)\bigr]\,dz.$
Since $Q^{\prime}(\theta_{b}+z)-Q^{\prime}(\theta_{b}-z)=\frac{d}{dz}\bigl(Q(\theta_{b}+z)+Q(\theta_{b}-z)\bigr)$, an integration by parts yields
$\partial_{\theta}u_{0}(r,\theta_{b})=\frac{k_{0}r}{2\pi}\int_{\gamma}\sin z\,e^{ik_{0}r\cos z}\bigl[Q(\theta_{b}+z)+Q(\theta_{b}-z)\bigr]\,dz.$
@@
$\partial_{\theta}u_{1}(r,\theta_{b})=\frac{k_{1}r}{2\pi}\int_{\gamma}\sin z\,e^{ik_{1}r\cos z}\bigl[S(\theta_{b}+z)+S(\theta_{b}-z)\bigr]\,dz.$
Now change variables in the medium-1 integrals by $w=w_{b}(z)$, so that $e^{ik_{1}r\cos z}=e^{ik_{0}r\cos w}$ by (7). The change of variables gives $dz=dw/w_{b}^{\prime}(z)$, while differentiating (7) implies $\sin z=\sin w\,w_{b}^{\prime}(z)/\nu$. Using $k_{1}=\nu k_{0}$, we obtain
$u_{1}(r,\theta_{b})$ $=\frac{1}{2\pi i}\int e^{ik_{0}r\cos w}\,\frac{S(\theta_{b}+z)-S(\theta_{b}-z)}{w_{b}^{\prime}(z)}\,dw,$
$\partial_{\theta}u_{1}(r,\theta_{b})$ $=\frac{k_{0}r}{2\pi}\int\sin w\,e^{ik_{0}r\cos w}\bigl[S(\theta_{b}+z)+S(\theta_{b}-z)\bigr]\,dw,$
with $z=z(w)$ the inverse map. Comparing with the corresponding expressions for $u_{0}$ and $\partial_{\theta}u_{0}$, and using Lemma 1.3, we obtain the two scalar identities
@@
$Q_{\mathrm{inc}}(\zeta)=\frac{1}{\zeta-(\theta_{i}+i\varepsilon)},\qquad Q(\zeta)=Q_{\mathrm{inc}}(\zeta)+Q_{\mathrm{scat}}(\zeta),$
@@
    $Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\bigr]+R(u),$
@@
    (Incident analyticity.) The scattered part is analytic at the incident spectral point $\zeta=\theta_{i}+i\varepsilon$ (Theorem 11.1).
@@
    (Far-field coefficient.) The diffracted far-field coefficient in medium 0 is
    $D(\theta,\theta_{i})=e^{-i3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\,Q_{\mathrm{scat}}(\theta),$
@@ 2.1 Lemniscatic curve
We work on the physical plus component $\Omega_{+}^{\mathrm{phys}}\subset\Sigma_{\mathrm{lem}}$ characterized by
@@
Origin of the Snell surface in the impedance-matched case. The Sommerfeld representations in the two media use the plane-wave factors $e^{ik_{0}r\cos z}$ (exterior) and $e^{ik_{1}r\cos z}$ (interior). On a wedge face $\theta=\theta_{b}$ one matches these factors by the analytic change of variables $z\mapsto w=w_{b}(z)$ determined implicitly by
$k_{0}\cos w_{b}(z)=k_{1}\cos z=\nu k_{0}\cos z.$
Thus the Snell map $w=w_{b}(z)$ satisfies $\cos w=\nu\cos z$ and is fixed by the radiation/limiting-absorption branch condition $w(z)\sim z+i\log\nu$ as $\Im z\to+\infty$. Writing $t=e^{iz}$ and $s=e^{iw}$, the identity $\cos w=\nu\cos z$ becomes
$\frac{1}{2}\Bigl(s+\frac{1}{s}\Bigr)=\nu\,\frac{1}{2}\Bigl(t+\frac{1}{t}\Bigr),\qquad\text{i.e.}\qquad s^{2}-\nu\Bigl(t+\frac{1}{t}\Bigr)s+1=0.$
@@
and a short calculation shows that the quadratic relation above is equivalent to $Y^{2}=2(t^{4}+1)$. The physical component $\Omega_{+}^{\mathrm{phys}}$ corresponds to the branch $|t|<1$ with $Y\to-\sqrt{2}$ as $t\to 0$, for which $s\sim t/\sqrt{2}$.
Quarter-period symmetry and orbit branches. The lemniscatic curve admits the order-four automorphism
$\tau:\Sigma_{\mathrm{lem}}\to\Sigma_{\mathrm{lem}},\qquad\tau(t,Y)=(it,-Y),$ (11)
which preserves $\Omega_{+}^{\mathrm{phys}}$. Let $w=w(t,Y)$ denote the (multi-valued) analytic function on $\Sigma_{\mathrm{lem}}$ defined by $e^{iw}=s(t,Y)$, with the physical branch fixed by $w(z)\sim z+i\log\nu$ as $\Im z\to+\infty$ (equivalently $s\sim t/\nu$ as $t\to 0$). Following the standard orbit construction for a right-angle wedge, we introduce four orbit branches $w_{m}$ by
$e^{iw_{m}(t,Y)}:=\bigl(s(\tau^{m}(t,Y))\bigr)^{(-1)^{m}},\qquad m\in\{0,1,2,3\}.$ (12)
@@ 2.2 Physical Snell exponential and spectral map
Differentiating $\cos w=\nu\cos z$ with $e^{iw}=s(t,Y)$ and $t=e^{iz}$ yields the algebraic derivative
$g^{\prime}(t,Y)=\frac{dw}{dz}=\frac{\sqrt{2}(t^{2}-1)}{Y}.$ (14)
@@
$s_{\zeta}:=\exp\bigl(i(\zeta-\theta_{w})\bigr).$ (15)
For $\zeta$ in the Sommerfeld strip (with $\varepsilon>0$ fixed), the physical branch map $\zeta\mapsto(t(\zeta),Y(\zeta))\in\Omega_{+}^{\mathrm{phys}}$ is defined by solving
@@
3 Weierstrass uniformization for $\tau=i$
We take the square lattice $\tau=i$ with Weierstrass invariants

### v3

# A Non-Reciprocal Elliptic Spectral Solution of the Right-Angle Penetrable Wedge Transmission Problem for $\nu=\sqrt{2}$

@@ Abstract
MSC 2020:
35J05, 78A45, 33E05, 45E10.
Keywords:
penetrable wedge diffraction; impedance-matched penetrable wedge; Sommerfeld integrals; Wiener–Hopf; Riemann–Hilbert; elliptic functions; lemniscatic curve; Weierstrass zeta-function.
@@
This manuscript was prepared with extensive assistance from the ChatGPT-5.2 Pro large language model. The AI system generated the source, including the exposition, symbolic derivations, and formula manipulations. The author reviewed, edited, and approved the final manuscript and accepts responsibility for its scientific content and any remaining errors.
@@
    §11 proves that $Q_{\mathrm{scat}}$ is analytic at the incident spectral point $\zeta=\zeta_{i}$ (limiting absorption), by the explicit exclusion of the incident label from the scattered pole set (and the injectivity of the injective uniformization).
@@
  • The complex spectral variable is $\zeta\in\mathbb{C}$ (Sommerfeld strip). The incident spectral pole is $\zeta_{i}=\theta_{i}+\mathrm{i}\varepsilon$ with $\varepsilon>0$ (limiting absorption), and the physical limit is $\varepsilon\to 0^{+}$.
  • We use the Sommerfeld integration parameter $z$ and the associated variable $t=e^{\mathrm{i}z}$.
  • The lemniscatic curve is $\Sigma_{\mathrm{lem}}:=\{(t,Y):Y^{2}=2(t^{4}+1)\}$. The physical sheet $\Omega_{\mathrm{phys}}^{+}\subset\Sigma_{\mathrm{lem}}$ is characterized by $|t|<1$ and $Y\to-\sqrt{2}$ as $t\to 0$.
  • $\zeta_{W}(u)$, $\wp(u)$, and $\wp^{\prime}(u)$ denote the Weierstrass zeta and elliptic functions with invariants $(g_{2},g_{3})=(4,0)$ (square lattice $\tau=\mathrm{i}$); the subscript distinguishes the Weierstrass zeta function from the spectral variable $\zeta$.
@@ 1 Introduction and setup
Canonical diffraction by angular regions originates with Sommerfeld’s exact half-plane solution and its Sommerfeld-integral representation [1], and its subsequent extension to wedge boundaries by functional-equation and factorization methods (notably the Malyuzhinets technique) [2, 5]. For penetrable (transmission) wedges the spectral reductions typically lead to generalized Wiener–Hopf or matrix factorization problems (see, e.g., [4, 7]) that do not admit closed forms in full generality (see also [6, 8, 9]). For a right-angled penetrable wedge formulation and analytical developments in certain parameter regimes, see Antipov and Silvestrov [10], Nethercote, Assier and Abrahams [11], and (in the no-contrast case) Kunz and Assier [12]. For high-frequency numerical-asymptotic methods for scattering by penetrable convex polygons—where local corner diffraction plays a central role—see Groth, Hewett and Langdon [13].
@@
Scope.
The analysis and the resulting closed form apply only to the special configuration $(\theta_{w},\nu,\rho)=(\pi/4,\sqrt{2},1)$. We do not claim an explicit closed-form solution for general penetrable wedges (arbitrary contrast and wedge angle), for which the standard spectral reductions lead to matrix/generalized Wiener–Hopf or multi-variable boundary-value problems; see [4, 7, 8, 10, 11, 12, 13].
Main results and where to find them.
@@
$(\Delta+k_{0}^{2})u_{0}=0\ \text{in }\{|\theta|>\theta_{w}\},\qquad(\Delta+k_{1}^{2})u_{1}=0\ \text{in }\{|\theta|<\theta_{w}\},$ (1)
@@
$u_{\mathrm{inc}}(r,\theta)=\exp\!\big(\mathrm{i}k_{0}r\cos(\theta-\theta_{i})\big),$ (2)
@@
$u_{0}(r,\pm\theta_{w})=u_{1}(r,\pm\theta_{w}),\qquad\partial_{\theta}u_{0}(r,\pm\theta_{w})=\partial_{\theta}u_{1}(r,\pm\theta_{w}),\qquad r>0.$ (3)
@@
$\zeta_{i}:=\theta_{i}+\mathrm{i}\varepsilon,\qquad\varepsilon>0.$ (4)
@@ 1.1 Sommerfeld representation and spectral split
$u^{(0)}(r,\theta)=\frac{1}{2\pi\mathrm{i}}\int_{\gamma}e^{\mathrm{i}k_{0}r\cos z}\,\big(Q(\theta+z)-Q(\theta-z)\big)\,\,\mathrm{d}z,$ (5)
@@
$u^{(1)}(r,\theta)=\frac{1}{2\pi\mathrm{i}}\int_{\gamma}e^{\mathrm{i}k_{1}r\cos z}\,\big(S(\theta+z)-S(\theta-z)\big)\,\,\mathrm{d}z,\qquad|\theta|<\theta_{w},$ (6)
@@ 1.3 Transmission conditions in spectral form
with the branch determined by $w_{b}(z)\sim z+\mathrm{i}\log\nu$ as $\Im z\to+\infty$. Write $w_{b}^{\prime}(z)=\,\mathrm{d}w_{b}/\,\mathrm{d}z$.
@@
$U(r,\theta):=\frac{1}{2\pi\mathrm{i}}\int_{\gamma}e^{\mathrm{i}kr\cos\zeta}\,H(\zeta)\,\,\mathrm{d}\zeta.$
@@
Assume $Q$ and $S$ are analytic in a common Sommerfeld strip and have sufficient decay so that integration by parts in $z$ produces no boundary terms. Fix a face $\theta=\theta_{b}$ and let $w=w_{b}(z)$ be as in (10). Then the impedance-matched transmission conditions (3) are equivalent to the pointwise spectral relation
$\begin{pmatrix}S(\theta_{b}+z)\\[2.0pt] S(\theta_{b}-z)\end{pmatrix}=\frac{1}{2}\begin{pmatrix}1+w_{b}^{\prime}(z)&1-w_{b}^{\prime}(z)\\ 1-w_{b}^{\prime}(z)&1+w_{b}^{\prime}(z)\end{pmatrix}\begin{pmatrix}Q(\theta_{b}+w_{b}(z))\\[2.0pt] Q(\theta_{b}-w_{b}(z))\end{pmatrix}$. (11)
@@
Evaluate (5) and its $\theta$-derivative at $\theta=\theta_{b}$. Differentiating under the integral sign gives
$\partial_{\theta}u_{0}(r,\theta_{b})=\frac{1}{2\pi\mathrm{i}}\int_{\gamma}e^{\mathrm{i}k_{0}r\cos z}\,\big(Q^{\prime}(\theta_{b}+z)-Q^{\prime}(\theta_{b}-z)\big)\,\,\mathrm{d}z.$
Since $Q^{\prime}(\theta_{b}+z)-Q^{\prime}(\theta_{b}-z)=\frac{\,\mathrm{d}}{\,\mathrm{d}z}\big(Q(\theta_{b}+z)+Q(\theta_{b}-z)\big)$, an integration by parts yields
$\partial_{\theta}u_{0}(r,\theta_{b})=\frac{k_{0}r}{2\pi}\int_{\gamma}\sin z\,e^{\mathrm{i}k_{0}r\cos z}\,\big(Q(\theta_{b}+z)+Q(\theta_{b}-z)\big)\,\,\mathrm{d}z.$
@@
$\partial_{\theta}u_{1}(r,\theta_{b})=\frac{k_{1}r}{2\pi}\int_{\gamma}\sin z\,e^{\mathrm{i}k_{1}r\cos z}\,\big(S(\theta_{b}+z)+S(\theta_{b}-z)\big)\,\,\mathrm{d}z.$
Now change variables in the medium-1 integrals by $w=w_{b}(z)$, so that $e^{\mathrm{i}k_{1}r\cos z}=e^{\mathrm{i}k_{0}r\cos w}$ by (10). The change of variables gives $\mathrm{d}z=\,\mathrm{d}w/w_{b}^{\prime}(z)$, while differentiating (10) implies $\sin z=\sin w\,w_{b}^{\prime}(z)/\nu$. Using $k_{1}=\nu k_{0}$, we obtain
$u_{1}(r,\theta_{b})=\frac{1}{2\pi\mathrm{i}}\int e^{\mathrm{i}k_{0}r\cos w}\,\frac{S(\theta_{b}+z)-S(\theta_{b}-z)}{w_{b}^{\prime}(z)}\,\,\mathrm{d}w,$
$\partial_{\theta}u_{1}(r,\theta_{b})=\frac{k_{0}r}{2\pi}\int\sin w\,e^{\mathrm{i}k_{0}r\cos w}\,\big(S(\theta_{b}+z)+S(\theta_{b}-z)\big)\,\,\mathrm{d}w,$
with $z=z(w)$ the inverse map. Comparing with the corresponding expressions for $u_{0}$ and $\partial_{\theta}u_{0}$, and using Lemma 1.3, we obtain
@@
$Q_{\mathrm{inc}}(\zeta)=\frac{1}{\zeta-(\theta_{i}+\mathrm{i}\varepsilon)},\qquad Q(\zeta)=Q_{\mathrm{inc}}(\zeta)+Q_{\mathrm{scat}}(\zeta),$
@@
    $Q_{\mathrm{scat}}(u)=\sum_{\ell\in I_{\mathrm{scat}}}C_{\ell}\big[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\big]+R(u),$
@@
    (Incident analyticity.) The scattered part is analytic at the incident spectral point $\zeta=\theta_{i}+\mathrm{i}\varepsilon$ (Theorem 11.1).
@@
    (Formal far-field coefficient.) A far-field coefficient in medium 0 obtained by steepest descent of the Sommerfeld integral is
    $D(\theta,\theta_{i})=e^{-\mathrm{i}3\pi/4}\sqrt{\frac{2}{\pi k_{0}}}\,Q_{\mathrm{scat}}(\theta),$
@@
Remark 1.6 (Reciprocity status).
For real transmission parameters, the physical penetrable-wedge scattering problem is expected to satisfy a reciprocity symmetry in the far field. The closed form constructed here is a meromorphic solution of the derived spectral functional system in the lemniscatic configuration; however, numerical tests of the far-field coefficient extracted in Theorem 13.1 indicate that the resulting coefficient is generally non-reciprocal. Accordingly, this manuscript presents an explicit elliptic solution of the spectral system and a corresponding formal far-field coefficient, but does not claim physical reciprocity of the diffraction coefficient.
@@ 2.1 Lemniscatic curve
We work on the physical plus component $\Omega_{\mathrm{phys}}^{+}\subset\Sigma_{\mathrm{lem}}$ characterized by
@@
Origin of the Snell surface in the impedance-matched case.
On a wedge face $\theta=\theta_{b}$ one matches the factors $e^{\mathrm{i}k_{0}r\cos z}$ and $e^{\mathrm{i}k_{1}r\cos z}$ by the analytic change of variables $z\mapsto w=w_{b}(z)$ determined implicitly by $\cos w=\nu\cos z$ and fixed by the radiation/limiting-absorption branch condition $w(z)\sim z+\mathrm{i}\log\nu$ as $\Im z\to+\infty$. Writing $t=e^{\mathrm{i}z}$ and $s=e^{\mathrm{i}w}$, the identity $\cos w=\nu\cos z$ becomes
$\frac{1}{2}\left(s+\frac{1}{s}\right)=\nu\,\frac{1}{2}\left(t+\frac{1}{t}\right),\qquad\text{i.e.}\qquad s^{2}-\nu\left(t+\frac{1}{t}\right)s+1=0.$
@@
and a short calculation shows that the quadratic relation above is equivalent to (12). The physical component $\Omega_{\mathrm{phys}}^{+}$ corresponds to the branch $|t|<1$ with $Y\to-\sqrt{2}$ as $t\to 0$, for which $s\sim t/\sqrt{2}$.
Quarter-period symmetry and orbit branches.
The lemniscatic curve admits the order-four automorphism
$\tau:\Sigma_{\mathrm{lem}}\to\Sigma_{\mathrm{lem}},\qquad\tau(t,Y)=(\mathrm{i}t,-Y),$ (14)
which preserves $\Omega_{\mathrm{phys}}^{+}$. Let $w=w(t,Y)$ denote the (multi-valued) analytic function on $\Sigma_{\mathrm{lem}}$ defined by $e^{\mathrm{i}w}=s(t,Y)$, with the physical branch fixed by $w(z)\sim z+\mathrm{i}\log\nu$ as $\Im z\to+\infty$ (equivalently $s\sim t/\nu$ as $t\to 0$). Following the standard orbit construction for a right-angle wedge, we introduce four orbit branches $w_{m}$ by
$e^{\mathrm{i}w_{m}(t,Y)}:=\big(s(\tau^{m}(t,Y))\big)^{(-1)^{m}},\qquad m\in\{0,1,2,3\}.$ (15)
@@ 2.2 Physical Snell exponential and spectral map
Differentiating $\cos w=\nu\cos z$ with $e^{\mathrm{i}w}=s(t,Y)$ and $t=e^{\mathrm{i}z}$ yields the algebraic derivative
$g^{\prime}(t,Y)=\frac{\,\mathrm{d}w}{\,\mathrm{d}z}=\frac{\sqrt{2}(t^{2}-1)}{Y},$ (17)
@@
$s_{\zeta}:=\exp\!\big(\mathrm{i}(\zeta-\theta_{w})\big).$ (18)
For $\zeta$ in the Sommerfeld strip (with $\varepsilon>0$ fixed), the physical branch map $\zeta\mapsto(t(\zeta),Y(\zeta))\in\Omega_{\mathrm{phys}}^{+}$ is defined by solving
@@
3 Weierstrass uniformization for the square lattice
We take the square lattice $\tau=\mathrm{i}$ with Weierstrass invariants